OpenStudy (anonymous):

1.Measure your own height and arm span in inches. You will likely need some help from a parent, guardian, or sibling to get accurate measurements. Record your measurements on the “Data Record” document. You will not need to submit the “Data Record” document as part of this project. Instead, you will use the record to help you complete part two. 2.Measure five additional people you know and record their arm spans and heights in inches

OpenStudy (anonymous):

2.Measure five additional people you know and record their arm spans and heights in inches. 3.Share the names and measurements of the five people you measured, as well as your own measurements, to your student partner in the class. You may request a method to getting a partner from your teacher. Together you and your partner will have 12 sets of data. Step Two: Creating Your Scatter Plot Using GeoGebra or a graphing software of your choice, draw your scatter plot and the line you think represents the line of best fit. Note: Directions for downloading and using GeoGebra can be found in the "Course Information area.” 1.Both you and your partner need to create a scatter plot of the 12 data points you both collected. 2.Draw the line that you feel best fits the data. 3.Copy and paste your scatter plot into a word processing document. 4.Share and discuss your scatter plot and the line of best fit with your partner. If your scatter plots or lines of best fit are different, discuss reasons why they might be different.

OpenStudy (anonymous):

Include your scatter plot and the answers to the following questions in your word processing document and submit to your instructor. Be sure to review how to save your files before getting started. You and your partner are free to discuss this part of the project. However, each of you will be responsible for submitting your own attachment of this file. 1.Which two points did you use to draw the line of best fit? 2.Write the equation of the line passing through those two points using the point-slope formula y - y1 = m(x - x1). Show all of your work. Remember to find the slope of the line first. 3.What does the slope of the line represent within the context of your graph? 4.Using the equation that you found in question 2, approximately how tall is a person whose arm span is 66 inches? 5.According to your line of best fit, what is the arm span of a 74-inch-tall person?

OpenStudy (anonymous):

@cwhansen can you try to help me

OpenStudy (anonymous):

@Noodle1998 @angelita321 @AlejandroMorel help!!! @NoelGreco

OpenStudy (anonymous):

@thomaster help me pls

thomaster (thomaster):

Where do you need help with?

OpenStudy (anonymous):

1.Which two points did you use to draw the line of best fit? 2.Write the equation of the line passing through those two points using the point-slope formula y - y1 = m(x - x1). Show all of your work. Remember to find the slope of the line first. 3.What does the slope of the line represent within the context of your graph? 4.Using the equation that you found in question 2, approximately how tall is a person whose arm span is 66 inches? 5.According to your line of best fit, what is the arm span of a 74-inch-tall person?

OpenStudy (anonymous):

That part

thomaster (thomaster):

This looks like an assignment, how can we help you with that?

OpenStudy (anonymous):

well i really dont understand the best of fit line

thomaster (thomaster):

Did you measure those 5 people and yourself?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Do u want me to send u the file

thomaster (thomaster):

Maybe your partner knows how to do this? or are you both having trouble with it?

OpenStudy (anonymous):

Well i couldnt find a partner so i did it myself so in total i measured 12 people

thomaster (thomaster):

Okay so you already created the scatterplot?

OpenStudy (anonymous):

yes

thomaster (thomaster):

Oke that's a very nice line. I think the first question isn't very hard,, which 2 points you take, that doesn't really matter since it's almost a perfect line

thomaster (thomaster):

You're having trouble writing the equation?

OpenStudy (anonymous):

yes. and the other questions after that

thomaster (thomaster):

Oke the equation will be in the form of y=ax+b b is the interception of the y axis but there is none since the line is going throught the origin. So it will be y=ax a is the slope to calculate the slope, you take 2 points and put them in this equation: \(\Large a=\dfrac{Y_2-Y_1}{X_2-X_1}\)

OpenStudy (anonymous):

ok so the points i choose are 73,73 and 70,70 so the slope would be 1 correct?

thomaster (thomaster):

Yes So the equation would be y=x

OpenStudy (anonymous):

so for y would substitute 73?

OpenStudy (anonymous):

@thomaster ?

thomaster (thomaster):

for what part of the question?

OpenStudy (anonymous):

i dont know i am really confused. im on this part>>Write the equation of the line passing through those two points using the point-slope formula y - y1 = m(x - x1). Show all of your work. Remember to find the slope of the line first.

thomaster (thomaster):

We just did that.. the equation is Y = x So for a height of 73 inch, the spam of the arms is 73 inch

OpenStudy (anonymous):

oh ok sorry. so now question 3

thomaster (thomaster):

The slope represents the relationship between the arm span and the height.

thomaster (thomaster):

and since the slope is 1, the lenght is equal to the arm span

OpenStudy (anonymous):

oh ok so then in question four the answer would be 66?

OpenStudy (anonymous):

and number 5 74?

OpenStudy (anonymous):

@thomaster

thomaster (thomaster):

yep :)

OpenStudy (anonymous):

ok thank you so much! and i will tag you in another question if i need your help